# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 5.2.1. SLMs for Healthcare
Hippocrates (Acikgoz et al., 2024) is an open-source framework designed for the medical domain, offering unrestricted access to its training data, codebase, checkpoints, and evaluation protocols  [^10] https://cyberiada.github.io/Hippocrates/. It integrates specialized medical knowledge through an extensive pre-training corpus from three datasets: Medical Guidelines, PMC-Patients (Zhao et al., 2022), and PubMedQA-contexts (Jin et al., 2019). This corpus, with around 300 million training tokens, ensures the model’s proficiency in medical terminology and practices. The Hippo series, a customized 7B model, undergoes continuous pre-training, instruction tuning, and reinforcement learning from human and AI feedback. It has specific datasets for instruction tuning and preference learning. Fine-tuned on Mistral and LLaMA2, Hippo surpasses existing 7B and 13B models, matching or exceeding 70B models in some cases. For example, Hippo-Mistral 7B can achieve 59.9% accuracy on MedQA, outperforming Meditron 70B (Chen et al., 2023a) whose performance is 58.5%. It is evaluated on six medical downstream tasks including MedMCQA (Pal et al., 2022), PubmedQA (Jin et al., 2019), MedQA (Jin et al., 2021), and the USMLE-step1, USMLE-step2, and USMLE-step3 exams (Jin et al., 2021).
BioMedLM (Bolton et al., 2024b) introduces a 2.7 billion parameter GPT-style autoregressive model, trained exclusively on PubMed abstracts and full articles (Gao et al., 2020). After fine-tuning, BioMedLM achieves robust biomedical question-answering performance, comparable with larger models, scoring 57.3% on MedMCQA (dev) and 69.0% on the MMLU medical genetics exam. It is also fine-tuned to provide useful answers to patient queries, demonstrating the potential of smaller models as a transparent, privacy-conserving, cost-effective, and environmentally-friendly foundation for specific NLP applications like biomedicine. The model is available on Hugging Face Hub  [^11] https://huggingface.co/stanford-crfm/BioMedLM.
AdaLM (Yao et al., 2021) develops small, fast, and effective domain-specific pre-trained language models. This approach continues the training of a medical domain-specific SLM atop an existing pre-trained model. To address domain distribution shifts, four combined strategies of domain adaptation and compression were tested: direct domain-specific pretraining, distillation followed by adaptation, adaptation followed by distillation, and simultaneous adaptation of both large and small models during initial distillation (adapt-ant-distill). For biomedical applications, a 16GB corpus from PubMed  [^12] https://pubmed.ncbi.nlm.nih.gov/ abstracts was used to adapt a BERT_base model (12 layers, 768 hidden size) (Devlin et al., 2019). The model was evaluated on three downstream tasks: JNLPBA (Collier et al., 2004), EBM PICO (Nye et al., 2018), and ChemProt (Krallinger et al., 2017), demonstrating that adaptation and distillation are optimal for developing task-agnostic, domain-specific small models.
MentalLLaMA (Yang et al., 2024f) models explainable mental health analysis as a text generation task and establishes the first IMHI dataset with 105K samples for instruction tuning of LLMs. It compiles data covering eight tasks from ten sources and uses expert-designed prompts with ChatGPT for explanations. The data undergo rigorous automatic and manual evaluations for accuracy and quality. Based on LLaMA2, MentaLLaMA includes a 7B variant and is the first open-source LLM for explainable mental health analysis on social media. Evaluated on the IMHI benchmark, MentaLLaMA matches state-of-the-art methods in accuracy and produces human-level explanations, showing strong generalizability on unseen tasks. The project code is available at  [^13] https://github.com/SteveKGYang/MentaLLaMA.
@@ 5.2.2. SLMs for Science
SciGLM (Zhang et al., 2024d) is a scientific language model capable of collegiate-level scientific reasoning. Central to this method is a novel self-reflective instruction annotation framework addressing data scarcity in science. This framework uses GPT-4 (Achiam et al., 2023) to generate step-by-step reasoning for unlabeled scientific problems, followed by self-reflective critique and revision. Self-reflective critique is a method for annotating scientific questions that involve three steps with designed prompts in Table 12: (i) employing a Chain of Thought (CoT) prompt, (see Prompt 1 in the table), to generate step-by-step answers; (ii) using a reflective prompt, (see Prompt 2), for responses with incorrect solutions, aiding GPT-4 in generating correct answers by analyzing its previous errors; and (iii) incorporating the actual answer directly into the prompt (see Prompt 3) to further assist in resolving the question. To eliminate low-quality questions and answers resulting from incomplete OCR, an instruction-quality classifier is trained and applied for further revisions. Leveraging this, the SciInstruct dataset is curated, encompassing physics, chemistry, mathematics, and formal proofs, enhancing the scientific and mathematical reasoning abilities of the ChatGLM-6B (Du et al., 2022) language model. Notably, SciGLM enhances the performance of the base model (ChatGLM3-6B-Base) by 3.06% in average scientific QA accuracy across CEval-Hard (Huang et al., 2024a), CEval-Sci (Huang et al., 2024a), MMLU-Sci (Hendrycks et al., 2021), SciEval (Sun et al., 2024a), and SciBench (Wang et al., 2024c). For broader community benefit, SciInstruct, SciGLM, the self-reflective framework, and fine-tuning code are available at  [^14] https://github.com/THUDM/SciGLM.
@@
Llemma (Azerbayev et al., 2023), an SLM tailored for mathematical reasoning, evolves from the Code Llama framework (Roziere et al., 2023). Trained using a standard auto-regressive language modeling objective, Llemma’s 7B model processes 200B tokens, including a 55B-token subset from the proposed Proof-Pile-2 dataset. This dataset comprises a rich mix of scientific papers, web data on mathematics, and mathematical code, with a content cutoff in April 2023. By continual pre-training on Proof-Pile-2, Llemma significantly enhances its few-shot performance across key mathematical benchmarks: MATH (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), OCWCourses (Lewkowycz et al., 2022), MMLU-STEM (Hendrycks et al., 2021), and SAT. This continual training enables Llemma to outperform all comparable open-weight language models. For instance, it achieves a 71.9% accuracy on the SAT, significantly higher than Code Llama’s 40.6%, underscoring the profound impact of targeted continual pre-training in boosting mathematical problem-solving capabilities.
ChemLLM (Zhang et al., 2024e) introduces a language model framework for chemistry. It incorporates ChemData, a dataset crafted for instruction tuning. This dataset transforms structured chemical data into dialogue formats suitable for training. Additionally, it includes ChemBench, a robust benchmark consisting of 4,100 multiple-choice questions. These questions span nine fundamental chemical tasks and aim to minimize linguistic style bias in model outputs. It is trained on the InternLM2-Base-7B model (Cai et al., 2024), initially enhancing general language ability with a multi-corpus of 1.7 million Q&A pairs from Hugging Face. In the second phase, it is fine-tuned using a mix of ChemData for chemical knowledge and the multi-corpus to retain general functionality. As a result, ChemLLM excels in various interdisciplinary chemical tasks through fluid dialog interactions, achieving core chemical task results comparable to GPT-4 (Achiam et al., 2023) and showing competitive performance with similar-sized LLMs. For instance, in the Mol2caption task, ChemLLM achieves a score of 92.6, outperforming GPT-3.5 and slightly below the score attained by GPT-4 (Achiam et al., 2023).
AstroLLaMA (Nguyen et al., 2023) effectively addresses the gap in the performance of LLMs in specialized fields like academic astronomy by introducing a domain-specific model. This 7-billion-parameter model, fine-tuned on over 300,000 astronomy abstracts from arXiv available on Kaggle  [^15] https://www.kaggle.com/Cornell-University/ arxiv using LLaMA-2 (Touvron et al., 2023b), adheres to causal language modeling objectives from its pre-training phase. It achieves a perplexity 30% lower than Llama2, indicating substantial domain adaptability. Despite fewer parameters, AstroLLaMA surpasses standard foundational models in generating insightful and scientifically relevant outputs. With its significant potential for further fine-tuning, AstroLLaMA is publicly available  [^16] https://huggingface.co/universeTBD/astrollama, enhancing research in astronomy through tasks like automated paper summarization and conversational agent development.
@@ 5.2.3. SLMs for Finance and Law
MindLLM (Yang et al., 2023a) introduces a series of bilingual, lightweight language models, developed from scratch, with configurations of 1.3 billion and 3 billion parameters. It incorporates experiences from large model development, covering data construction, model architecture, evaluation, and applications. The primary English pre-training data is sourced from the Pile dataset (Gao et al., 2020), and the Chinese pre-training data combines resources from WuDao (Yuan et al., 2021), CBook  [^17] https://github.com/FudanNLPLAB/CBook-150K, and various Chinese websites. MindLLM consistently matches or surpasses other large models in public benchmarks. Bilingual training from scratch is advocated as a superior method for capacity learning and avoiding catastrophic forgetting, featuring a specialized instruction tuning framework designed to enhance smaller models’ capabilities effectively. Additionally, supervised fine-tuning in law and finance explores specific vertical domains. In law, publicly available legal data, scenario-based Q&A, legal references from LaW-GPT (Hongcheng Liu, 2023), and NLP-based legal tasks from DISC-LawLLM (Yue et al., 2023) are used. Due to the absence of public objective benchmarks in the legal field, the robust ChatGPT model serves as an evaluator for reasoning on manually constructed data. In finance, EastMoney  [^18] https://www.eastmoney.com/default.html is selected as the data source. Financial sentiment data is formatted into a text classification task, where inputs include financial news and outputs are labels for five categories: positive, negative, neutral, very negative, and very positive. MindLLM-1.3B and 3B models demonstrate superior performance, achieving accuracy rates of 47.79% and 46.40%, respectively, to similar-sized models and outperform ChatGLM2-6B (45.79%) and Open-LLaMA-7B (28.38%) after fine-tuning.
Insights: We draw several key insights from the development of domain-specific SLMs: Adapting SLMs to domain-specific data is a common practice for acquiring domain-specific SLMs, prompting many to create their datasets (Yang et al., 2024f; Nguyen et al., 2023; Zhang et al., 2024e, d). These datasets are often annotated using LLMs like GPT-4 and used to train or fine-tune general models such as LLaMa-2-7B (Acikgoz et al., 2024; Bolton et al., 2024b). To ensure the data quality, specialized annotation frameworks are developed, such as Zhang et al. (2024d). In domains with abundant corpora, training a general model from scratch and fine-tuning it using SFT (Yang et al., 2023a) is practical. Bilingual settings during training can prevent catastrophic forgetting. Distilling general capabilities from LLMs while integrating domain-specific knowledge from corpora is another method for developing domain-specific SLMs (Yao et al., 2021).
6. SLMs for LLMs
In this section, we provide a comprehensive review of how SLMs enhance LLMs. While LLMs are robust, they face challenges such as latency during inference, labor-intensive fine-tuning, noise filtration issues in retrieval, suboptimal zero-shot performance, copyright infringement risks, and evaluation difficulties. SLMs can help LLMs to alleviate these issues. Research in this field can be categorized into five primary areas: (i) using SLMs for reliable LLM generation; (ii) extracting prompts for LLMs using SLMs; (iii) fine-tuning LLMs with SLMs; (iv) applying SLMs in LLM applications; and (v) evaluating LLMs using SLMs. A summary of representative work in each category along with their key point is given in Table 13. Next, we introduce each category in detail.
@@ 6. SLMs for LLMs
POLAR (Zhao et al., 2023b) Uses a small BERT model to calibrate LLM responses and assign risk scores.
Hallucination Detector in NMT (Xu et al., 2023a) Uses lightweight classifiers to detect hallucinations in Neural Machine Translation.
SAPLMA (Azaria and Mitchell, 2023) Uses a BERT Small Language Model as a classifier to assess the truthfulness of statements accurately.
@@
SuperContext (Yang et al., 2024g) Specific SLM enhances ICL by providing confidence and predictions to overcome out-of-domain challenges.
@@
SKR (Wang et al., 2023c) Trains a small model to detect its self-knowledge for better use of external knowledge.
SlimPLM (Tan et al., 2024) Detects missing knowledge in LLMs with a slim proxy model, enhancing the LLM’s knowledge integration.
In-Context RALM (Ram et al., 2023) Trains a RoBERTa-based reranker for top-k BM25 documents using LM signals to enhance LM gains.
CRAG (Yan et al., 2024) Trains a lightweight evaluator to assess document quality and triggers actions based on confidence levels.
@@
Prompt Stealing Attacks (Sha and Zhang, 2024) Uses small models fine-tuned as parameter extractors to facilitate hierarchical prompt reconstruction.
Output2prompt (Zhang et al., 2024f) Uses a sparse encoder-decoder-based T5 small model to reverse-engineer LLM inputs from outputs.
Model Purifying (Li et al., 2024e) Uses SLMs to ensemble with LLMs, mitigating negative effects from uncurated data.
@@
Emulated Fine-tuning (Mitchell et al., 2024) Emulates pre-training and fine-tuning at different scales by summing base log probabilities with behavior deltas.
@@
SLM for LLM applications SLCoLM (Tang et al., 2024b) Uses SLM predictions to guide the LLM generation process in Chinese Entity Relation Extraction.
HEF (Yang et al., 2024d) Uses SLMs as flexible plugins to improve LLM’s nuanced understanding.
Contrastive decoding (Li et al., 2023d) Enhances text quality by maximizing the difference between expert and amateur log probabilities.
SLM for LLM evaluation SLIDE (Zhao et al., 2024d) Utilizes SLMs trained via contrast learning to distinguish and score responses in dialogue scenarios effectively.
@@ 6.1. SLM for Reliable LLM Generation
Enhancing Calibration and Hallucination Detection of LLMs As illustrated in Figure 18 (a), to calibrate LLMs, an SLM processes both questions and LLM-generated answers to predict calibrated confidence. This training involves minimizing the discrepancy between estimated calibration error and predicted confidence score. For instance, APRICOT (Ulmer et al., 2024) uses an auxiliary DeBERTaV3 model (He et al., 2023a) to assess LLM confidence in open-question scenarios, aiming to improve uncertainty expression and response adjustment. Similarly, Zhao et al. (2023b) has developed a self-supervised approach that generates risk scores for each response to calibrate LLM confidence, utilizing a small BERT model (Devlin et al., 2019) to synchronize LLM outputs with other weak supervision sources. As shown in Figure 18 (b), for hallucination detection, an SLM analyzes LLM internal states to output the likelihood of hallucination. This process uses supervised data obtained by testing the knowledge boundaries of the LLM. In neural machine translation, Xu et al. (2023a) develop a lightweight detector that analyzes token contributions to hallucinations, outperforming both model-free baselines and quality estimation classifiers. Furthermore, SAPLMA (Azaria and Mitchell, 2023) found that LLM internal states can signal the truthfulness of statements, with a small BERT classifier trained to differentiate correct from incorrect predictions achieving accuracies of 71% to 83%.
@@
Enhancing Retrieval-Augmented Generation Generally, as shown in Figure 19, SLMs can also serve as proxy models to evaluate the familiarity of LLMs with user queries, determining whether LLMs need to retrieve additional information or can respond directly. For example, SlimPLM (Tan et al., 2024) is a small proxy model that assesses the necessity for LLM retrieval by generating heuristic answers. High-quality responses indicate that LLMs can handle queries independently, whereas lower-quality outputs require further retrieval. Additionally, Self-Knowledge Guided Retrieval (SKR) (Wang et al., 2023c) enables SLMs to autonomously decide when LLMs should operate independently, based on their self-assessment of knowledge limitations. Further, SELF-RAG (Asai et al., 2024) improves the factual accuracy and quality of LLM outputs through on-demand retrieval and self-reflection. This method employs a small critic language model to issue reflective markers and make binary decisions regarding the need for further information retrieval. Moreover, some studies utilize SLMs to evaluate the relevance of retrieved documents. LongLLMLingua (Jiang et al., 2024) employs SLMs to calculate the relevance of documents to a query $x^{que}$ using perplexity, as formalized by the equation:
@@
where $x^{que}_{i}$ is the $i$-th token in the query sequence, $x_{k}^{doc}$ is the retrieved document, and $N_{c}$ is the total number of tokens in the query. $p_{\text{SLM}}$ represents the probability generated by an SLM. CRAG (Yan et al., 2024) employs SLMs as evaluators of document relevance in the same way. In addition, other research employs SLMs as re-rankers to refine the order of documents provided by initial retrieval efforts such as BM25 (Robertson et al., 2009). In-Context RALM (Ram et al., 2023) positions SLMs as rankers, optimizing the document sequence with a fine-tuning process on RoBERTa (Liu et al., 2019) as defined by the loss function:
@@
where $x_{\leq s_{i}}$ is a prefix sampled from the training data, $y=x_{s_{i}+1},\ldots,x_{s_{i}+s}$ represents the text to be generated in the next stride, $p_{\theta}(y|d_{i};x_{\leq si})$ denotes the probability of the LLM generating $y$ given $d_{i}$ and $x_{\leq si}$, and $p_{\text{rank}}(d_{i}|x_{\leq s_{j}})$ is the ranking score of $d_{i}$.
@@
Enhancing Reasoning Capabilities of LLMs As illustrated in Figure 20, SLMs enhance LLMs reasoning by transferring task knowledge to in-context examples, effectively reducing hallucinations. While In-context Learning (ICL) generally handles few-shot learning with 16 to 32 examples, it struggles when faced with extensive supervised data. SLMs, specialized in task-specific training, complement the broader domain knowledge of extensively pre-trained LLMs. For example, SuperICL (Xu et al., 2023c) incorporates SLMs as plugins for efficiently executing supervised tasks. It predicts labels for contextual examples and integrates these predictions with the input text and actual labels to enhance knowledge transfer, thereby boosting the understanding and responsiveness of LLMs. SuperContext (Yang et al., 2024g) tackles challenges that LLMs encounter with new tasks and out-of-distribution data in natural language understanding by synergizing SLM outputs with LLM prompts during inference. This integration merges model predictions with their confidence levels, effectively leveraging SLM task-specific knowledge and LLM domain expertise. Furthermore, SLMs efficiently decompose complex reasoning by breaking tasks into simpler components, as demonstrated in (Wu et al., 2024a). This strategy not only increases efficiency but also reduces deployment costs when SLMs and LLMs are used collaboratively, transforming complex tasks into manageable segments.
@@
Alleviate Copyright and Privacy Issues of LLMs As depicted in Figure 21, SLMs can assist LLMs in addressing copyright and privacy concerns arising from online data collection. By training on selectively curated data subsets, SLMs effectively reduce copyright infringement and privacy risks, although they are less effective than full-scale LLMs. To harness the combined benefits of both models, Li et al. (2024e) integrates untrusted LLMs with benign SLMs using the CP-$\Delta$ KL algorithm to mitigate adverse effects while preserving performance. The equation is:

### v2

@@ 5.2.1. SLMs for Healthcare
Hippocrates (Acikgoz et al., 2024) is an open-source medical language model framework with free access to its data, codebase, checkpoints, and protocols  [^7] https://cyberiada.github.io/Hippocrates/. It utilizes a medical pre-training corpus from Medical Guidelines, PMC-Patients (Zhao et al., 2022), and PubMedQA-contexts (Jin et al., 2019), totaling about 300M tokens. The Hippo series, a 7B model, undergoes continuous pre-training, instruction tuning, and RLHF. Fine-tuned on Mistral and Llama-2, it rivals 70B models in some evaluation. For example, Hippo-Mistral 7B scores 59.9% on MedQA, outperforming Meditron 70B (Chen et al., 2023a) at 58.5%. BioMedLM (Bolton et al., 2024b), a 2.7B GPT-style model trained on PubMed content (Gao et al., 2020), excels in biomedical QA after fine-tuning, achieving 57.3% on MedMCQA (dev) and 69.0% on MMLU medical genetics exams. Available on Hugging Face Hub  [^8] https://huggingface.co/stanford-crfm/BioMedLM. AdaLM (Yao et al., 2021) enhances domain-specific SLMs by continuing training on a medical-focused SLM atop a general pre-trained model. It emperically validates adaptation then distillation is the most effective distillation way. AdaLM modified a BERT_base model (12 layers, 768 hidden size) (Devlin et al., 2019) with a 16GB PubMed  [^9] https://pubmed.ncbi.nlm.nih.gov/ abstracts corpus. MentalLLaMA (Yang et al., 2024i) introduces the first IMHI dataset for mental health analysis and the first open-source LM for explainable analysis on social media. The IMHI is compiled from ten sources, totaling 105K samples. Expert-designed mental health analysis prompts are employed via ChatGPT for explanations. Based on Llama-2-7B, MentalLLaMA is instruction-tuned on this data and matches top methods in accuracy on the IMHI test set. Project code is available at  [^10] https://github.com/SteveKGYang/MentaLLaMA.
@@ 5.2.2. SLMs for Science
SciGLM (Zhang et al., 2024d) is a collegiate-level scientific language model overcoming data scarcity with a self-reflective instruction annotation framework. Utilizing GPT-4 (Achiam et al., 2023), it generates detailed reasoning for unlabeled scientific problems through three steps with designed prompts in Table 11: (i) CoT prompt for step-by-step answers (Prompt 1), (ii) reflective prompt for correcting errors (Prompt 2), and (iii) integrating the correct answer for clarity (Prompt 3). The SciInstruct dataset spans physics, chemistry, math, and proofs, tuning ChatGLM-6B’s (Du et al., 2022) reasoning abilities. SciGLM boosts the base model’s (ChatGLM3-6B-Base) scientific QA accuracy by 3.06% on benchmarks such as CEval-Hard (Huang et al., 2024a), CEval-Sci (Huang et al., 2024a), MMLU-Sci (Hendrycks et al., 2021), SciEval (Sun et al., 2024a), and SciBench (Wang et al., 2024e). Llemma (Azerbayev et al., 2023), an SLM derived from CodeLlama (Roziere et al., 2023), specializes in mathematical reasoning. By continual pre-training, its 7B model is evolved on 55B tokens from the newly created Proof-Pile-2 dataset, which includes scientific papers, math web content, and mathematical code up until April 2023, to enhance few-shot capabilities. It excels in mathematical benchmarks like MATH (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), OCWCourses (Lewkowycz et al., 2022), MMLU-STEM (Hendrycks et al., 2021), and SAT, surpassing all comparable open-weight models. ChemLLM (Zhang et al., 2024e) is a chemistry-focused language model that utilizes its proposed ChemData, a dataset designed for instruction tuning that transforms chemical data into dialogue format for training. ChemLLM is based on InternLM2-Base-7B (Cai et al., 2024a), initially enhancing its language skills with a multi-corpus of 1.7 million Q&A pairs from Hugging Face, then fine-tunes using ChemData and the multi-corpus to maintain its general capabilities. ChemLLM excels in interdisciplinary chemical tasks within the proposed ChemBench, achieving results comparable to GPT-4 (Achiam et al., 2023) and outperforming GPT-3.5 with a score of 92.6 in Mol2caption, slightly below that of GPT-4. AstroLLaMA (Nguyen et al., 2023) introduces an astronomy-focused language model. Based on Llama-2-7B (Touvron et al., 2023b) and enhanced via continual pre-training, it has been developed using over 300K astronomy abstracts from arXiv  [^11] https://www.kaggle.com/Cornell-University/arxiv. AstroLLaMA achieves 30% lower perplexity than Llama-2-7B, indicating substantial improvements in domain adaptability. AstroLLaMA is available  [^12] https://huggingface.co/universeTBD/astrollama for tasks such as automated paper summarization and conversational agent development in astronomy.
@@ 5.2.3. SLMs for Finance and Law
MindLLM (Yang et al., 2023a) introduces a bilingual (Chinese and English) SLM, pretrained on the Pile dataset (Gao et al., 2020) for English and WuDao (Yuan et al., 2021), CBook, and various Chinese web content for Chinese. Bilingual training enhances capacity and prevents catastrophic forgetting. It explores specific domains such as law and finance through supervised fine-tuning. In law, it utilizes publicly available legal data, scenario-based Q&A from LaW-GPT (Hongcheng Liu, 2023), and NLP-based legal tasks from DISC-LawLLM (Yue et al., 2023). In finance, EastMoney  [^13] https://www.eastmoney.com/default.html is selected as the data source.
Insights: We draw several key insights from the development of domain-specific SLMs: Adapting SLMs to domain-specific data is a common practice for acquiring domain-specific SLMs, prompting many to create their datasets (Yang et al., 2024i; Nguyen et al., 2023; Zhang et al., 2024e, d). These datasets are often annotated using LLMs like GPT-4 and used to continual pre-train or fine-tune general models such as LLaMa-2-7B (Acikgoz et al., 2024; Bolton et al., 2024b). To ensure the data quality, specialized annotation frameworks are developed, such as SciGLM (Zhang et al., 2024d). In domains with abundant corpora, training a general model from scratch and fine-tuning it using SFT (Yang et al., 2023a) is practical. Bilingual settings during training can prevent catastrophic forgetting. Distilling general capabilities from LLMs while integrating domain-specific knowledge from corpora is another method for developing domain-specific SLMs (Yao et al., 2021).
@@
POLAR (Zhao et al., 2023b) Using a BERT model to calibrate LLM responses.
Hallucination Detector in NMT (Xu et al., 2023a) Using lightweight classifiers to detect hallucinations in Neural Machine Translation.
SAPLMA (Azaria and Mitchell, 2023) Using a BERT Small Language Model as a classifier to assess the truthfulness of statements accurately.
@@
SuperContext (Yang et al., 2024j) Specific SLM enhances ICL by providing confidence and predictions to overcome out-of-domain challenges.
@@
SKR (Wang et al., 2023c) Training a small model to detect its self-knowledge for better use of external knowledge.
SlimPLM (Tan et al., 2024) Detecting missing knowledge in LLMs with a slim proxy model, enhancing the LLM’s knowledge integration.
In-Context RALM (Ram et al., 2023) Training a RoBERTa-based reranker for top-k BM25 documents using LM signals to enhance LM gains.
CRAG (Yan et al., 2024) Training a lightweight evaluator to assess document quality and trigger actions based on confidence levels.
GSR (Huang et al., 2024c) Training a Generative Sub-graph Retriever (GSR) for relation chain in RAG when retrieving from knowledge graphs.
@@
Prompt Stealing Attacks (Sha and Zhang, 2024) Using small models fine-tuned as parameter extractors to facilitate hierarchical prompt reconstruction.
Output2prompt (Zhang et al., 2024f) Using a sparse encoder-decoder-based T5 small model to reverse-engineer LLM inputs from outputs.
Model Purifying (Li et al., 2024f) Using SLMs to ensemble with LLMs, mitigating negative effects from uncurated data.
@@
Emulated Fine-tuning (Mitchell et al., 2024) Emulating pre-training and fine-tuning at different scales by summing base log probabilities with behavior deltas.
@@
Weak-to-Strong Search (Zhou et al., 2024) Framing LLM alignment as a test-time greedy search to maximize the log-probability difference between tuned and untuned SLMs.
SLM for LLM applications SLCoLM (Tang et al., 2024b) Using SLM predictions to guide the LLM generation process in Chinese Entity Relation Extraction.
HEF (Yang et al., 2024e) Using SLMs as plugins to improve LLM’s nuanced understanding.
Contrastive decoding (Li et al., 2023e) Enhancing text quality by maximizing the difference between expert and amateur log probabilities.
SLM for LLM safety Llama Guard (Inan et al., 2023) An LLM-based input-output safeguard model geared towards Human-AI conversation use cases.
SLM as Guardian (Kwon et al., 2024) A smaller LLM for both harmful query detection and safeguard response generation.
SLM for LLM evaluation SLIDE (Zhao et al., 2024d) Utilizing SLMs trained via contrast learning to distinguish and score responses in dialogue scenarios effectively.
Kuhn et al. (2023) An SLM is used as the natural language inference classifier.
SelfCheckGPT (Manakul et al., 2023) An SLM is used to calculate BERTScore.
Factscore (Min et al., 2023) An SLM functions as the natural language inference classifier.
6. SLMs for LLMs
In this section, we provide a comprehensive review of how SLMs enhance LLMs. While LLMs are robust, they face challenges such as latency during inference, labor-intensive fine-tuning, noise filtration issues in retrieval, suboptimal zero-shot performance, copyright infringement risks, and evaluation difficulties. SLMs can help LLMs to alleviate these issues. Research in this field can be categorized into five primary areas: (i) using SLMs for reliable LLM generation; (ii) extracting prompts for LLMs using SLMs; (iii) fine-tuning LLMs with SLMs; (iv) applying SLMs in LLM applications; (v) utilizing SLMs as guardian; and (vi) evaluating LLMs using SLMs. A summary of representative work in each category along with their key point is given in Table 12. Next, we introduce each category in detail.
@@ 6.1. SLM for Reliable LLM Generation
Enhancing Calibration and Hallucination Detection of LLMs As illustrated in Figure 20 (a), to calibrate LLMs, an SLM processes both questions and LLM-generated answers to predict calibrated confidence. This training involves minimizing the discrepancy between estimated calibration error and predicted confidence score. For instance, APRICOT (Ulmer et al., 2024) uses an auxiliary DeBERTaV3 model (He et al., 2023a) to assess LLM confidence in open-question scenarios, aiming to improve uncertainty expression and response adjustment. Similarly, POLAR (Zhao et al., 2023b) has developed a self-supervised approach that generates risk scores for each response to calibrate LLM confidence, utilizing a small BERT model (Devlin et al., 2019) to synchronize LLM outputs with other weak supervision sources. As shown in Figure 20 (b), for hallucination detection, an SLM analyzes LLM internal states to output the likelihood of hallucination. This process uses supervised data obtained by testing the knowledge boundaries of the LLM. In neural machine translation, Xu et al. (2023a) develop a lightweight detector that analyzes token contributions to hallucinations, outperforming both model-free baselines and quality estimation classifiers. Furthermore, SAPLMA (Azaria and Mitchell, 2023) found that LLM internal states can signal the truthfulness of statements, with a small BERT classifier trained to differentiate correct from incorrect predictions achieving accuracies of 71% to 83%.
@@
Enhancing Retrieval-Augmented Generation Generally, as shown in Figure 21, SLMs can also serve as proxy models to evaluate the familiarity of LLMs with user queries, determining whether LLMs need to retrieve additional information or can respond directly. For example, SlimPLM (Tan et al., 2024) is a small proxy model that assesses the necessity for LLM retrieval by generating heuristic answers. High-quality responses indicate that LLMs can handle queries independently, whereas lower-quality outputs require further retrieval. Additionally, Self-Knowledge Guided Retrieval (SKR) (Wang et al., 2023c) enables SLMs to autonomously decide when LLMs should operate independently, based on their self-assessment of knowledge limitations. Further, SELF-RAG (Asai et al., 2024) improves the factual accuracy and quality of LLM outputs through on-demand retrieval and self-reflection. This method employs a small critic language model to issue reflective markers and make binary decisions regarding the need for further information retrieval. Moreover, some studies utilize SLMs to evaluate the relevance of retrieved documents. LongLLMLingua (Jiang et al., 2024b) employs SLMs to calculate the relevance of documents to a query $x^{que}$ using perplexity, as formalized by the equation:
@@
where $x^{que}_{i}$ is the $i$-th token in the query sequence, $x_{k}^{doc}$ is the retrieved document, and $N_{c}$ is the total number of tokens in the query. $p_{\text{SLM}}$ represents the probability generated by an SLM. CRAG (Yan et al., 2024) employs SLMs as evaluators of document relevance in the same way. RA-ISF (Liu et al., 2024f) trains a small language model that checks the base LLM in self-knowledge, relevance judgment, and question decomposition. In addition, some research employs SLMs as re-rankers to refine the order of documents provided by initial retrieval efforts such as BM25 (Robertson et al., 2009). In-Context RALM (Ram et al., 2023) positions SLMs as rankers, optimizing the document sequence with a fine-tuning process on RoBERTa (Liu et al., 2019) as defined by the loss function:
@@
where $x_{\leq s_{i}}$ is a prefix sampled from the training data, $y=x_{s_{i}+1},\ldots,x_{s_{i}+s}$ represents the text to be generated in the next stride, $p_{\theta}(y|d_{i};x_{\leq si})$ denotes the probability of the LLM generating $y$ given $d_{i}$ and $x_{\leq si}$, and $p_{\text{rank}}(d_{i}|x_{\leq s_{j}})$ is the ranking score of $d_{i}$. Lastly, some studies leverage SLMs to retrieve sub-graphs when utilizing knowledge graphs as external sources. Huang et al. (2024c) introduces the Generative Sub-graph Retriever (GSR), which employs SLMs to predict relation chains for answering questions, offering a cost-effective alternative to training LLMs. Specifically, it uses customized T5 (220M, 770M, and 3B) (Raffel et al., 2020) as retrievers to enhance LLM readers, including Llama2-chat-7B (Touvron et al., 2023b) and Llama3-instruct-8B (Dubey et al., 2024), on the WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018) datasets.
@@
Enhancing Reasoning Capabilities of LLMs As illustrated in Figure 22, SLMs enhance LLMs reasoning by transferring task knowledge to in-context examples, effectively reducing hallucinations. While In-context Learning (ICL) generally handles few-shot learning with 16 to 32 examples, it struggles when faced with extensive supervised data. SLMs, specialized in task-specific training, complement the broader domain knowledge of extensively pre-trained LLMs. For example, SuperICL (Xu et al., 2023c) incorporates SLMs as plugins for efficiently executing supervised tasks. It predicts labels for contextual examples and integrates these predictions with the input text and actual labels to enhance knowledge transfer, thereby boosting the understanding and responsiveness of LLMs. SuperContext (Yang et al., 2024j) tackles challenges that LLMs encounter with new tasks and out-of-distribution data in natural language understanding by synergizing SLM outputs with LLM prompts during inference. This integration merges model predictions with their confidence levels, effectively leveraging SLM task-specific knowledge and LLM domain expertise. Furthermore, SLMs efficiently decompose complex reasoning by breaking tasks into simpler components, as demonstrated in (Wu et al., 2024a). This strategy increases efficiency and reduces deployment costs when SLMs and LLMs are used collaboratively, transforming complex tasks into manageable segments.
@@
Alleviate Copyright and Privacy Issues of LLMs LLMs pose significant security risks due to their tendency to memorize training data, leading to potential privacy breaches and copyright infringement. As depicted in Figure 23, SLMs can assist LLMs in addressing copyright and privacy concerns arising from online data collection. By training on selectively curated data subsets, SLMs effectively reduce copyright infringement and privacy risks, although they are less effective than full-scale LLMs. To harness the combined benefits of both models, Li et al. (2024f) integrates untrusted LLMs with benign SLMs using the CP-$\Delta$ KL algorithm to mitigate adverse effects while preserving performance. The equation is: